OpenStudy (anonymous):

-1/7x+4(2+2/7x)= -8 please show how to solve this multistep equation.

OpenStudy (anonymous):

EASSSYYYYYY - I will write the steps for you.

OpenStudy (anonymous):

x = -16

OpenStudy (anonymous):

can u show me the steps pls

OpenStudy (anonymous):

-1/7x+4(2+2/7x)= -8 First distribute the 4 -1/7x + 8 + 8/7x = -8 Then combine like terms -1/7x + 8/7x 8 + x = -8 Then subtract 8 from both sides. -8 -8 --------------- x = -16

OpenStudy (anonymous):

oooo wow thanks a bunch

OpenStudy (anonymous):

No problem!! let me know if you need anymore help!! :)

OpenStudy (anonymous):

i have a few more that are throwing me off, do u mind if i send them to u ?

OpenStudy (anonymous):

sure post it as new question ;)

OpenStudy (anonymous):

ok